• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 07:55:49 on June 5,2002

    Entry number 85273

    keyword=Owl Shift Summary

    0:00 Beam is off. MCC investigating why they cannot deliver stable 100uA to us. Problem is with injector.
    0:36 Beam returns. 100uA. Start run 1172. Sieve in Left, no collimator in Right.
    1:00 MCC reduces current to 80uA to stabilize beam delivery and to investigate injector problems. Many trips, quite inefficient data taking.
    1:10 Beam goes away for more injector work.
    1:25 Beam returns, 80uA, a bit better than before. Still many trips.
    1:50 Trip frequency increasing. MCC lowers current, still doesn't seem to help much.
    2:08 Beam gone again. MCC investigates why it is so terrible.
    2:27 Beam comes back, 80uA. More stable now.
    2:52 Beam deteriorates again.
    3:00 Stop data taking. Move in 15cm LH2 target.
    3:02 Target in place. Asking for beam.
    3:20 Asking for 50uA instead of 80uA in hopes of more stable beam.
    3:55 Changing left arm collimator from sieve to open.
    4:02 Collimator in place. Asking for 15uA.
    4:08 Start data taking, LH2 15cm, collimators on both arms open.
    4:23 Moving in 6msr collimator on left arm.
    4:28 Start data taking
    4:32 End run. MCC reports that North Linac dropped to Restricted Access, and many magnets have tripped. No beam for a while.
    6:30 MCC says "at least another 2 hours until beam". Oh well.